The current in a circuit with an external resistance of 3.75 ohm is 0.5 . when a resistance of 1ohm is introduce into the circuit the current becomes 0.4 .Then what will be the emf of the power source ?

Grade:12

2 Answers

Eshan
5 years ago
Let the EMF and internal resistance of the battery beEandr, respectively.

When external resistance of3.75\Omegais connected,

E=i(R+r)
\implies E=0.5(3.75+r)
When external resistance of1\Omegais added to the circuit,

E=i(R+R_1+r)
\implies E=0.4(3.75+1+r)=0.4(4.75+r)
E=2V
Gitanjali Rout
5 years ago
q] The current in a circuit with an external resistance of 3.75 ohm is 0.5 . when a resistance of 1ohm is introduce into the circuit the current becomes 0.4 .Then what will be the emf of the power source ?
E=2V\implies E=0.4(3.75+1+r)=0.4(4.75+r)E=i(R+R_1+r)is added to the circuit,1\OmegaWhen external resistance of\implies E=0.5(3.75+r)E=i(R+r)is connected,3.75\Omega, respectively.When external resistance ofrandELet the EMF and internal resistance of the battery be